OpenStudy (anonymous):

im stuck on this problem : Solve 3x^2 = –12x – 15

OpenStudy (across):

Let's simplify this; do we want to divide both sides by 3?

OpenStudy (anonymous):

yea

OpenStudy (across):

What would we get then? :)

OpenStudy (anonymous):

x^2 +6x-5

OpenStudy (across):

That's right:\[3x^2 = –12x – 15,\]\[x^2+4x+5=0.\]We moved everything to the left side. Now, do you know how to factor that?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):



OpenStudy (anonymous):

ok so any negative number is "i "

OpenStudy (across):

Very close!\[\frac{-4\pm\sqrt{16-20}}{2}=-2\pm i\](remember that \(\sqrt{-1}=i\)).

OpenStudy (anonymous):

thanks!